How many liters of a 20% soul utopian of a chemical should be mixed with 15l of a 50% solutionto get a 30% solution ?

Answer by stanbon(75887)

How many liters of a 20% solution of a chemical should be mixed with 15l of a 50% solution to get a 30% solution ?
-----------
Equation:
active + active = active
0.20x + 0.50*151 = 0.30(x+151)
----
20x + 50*151 = 30x + 30*151
-10x = -20*151
x = 151/2 = 75.5 liters (amt of 20% solution needed)
--------------
Cheers,
Stan H.
--------------